OpenStudy (anonymous):

Urgent! Please help me simplify 2/x^2/10/x^5.Thanks, I'm a bit rusty on my math.

OpenStudy (abb0t):

Just to clarify, is it: \(\large \sf \color{red}{\frac{2}{x^2}÷ \frac{10}{x^5}}\), yes?

OpenStudy (anonymous):

@abb0t yes, except it's written like this-

OpenStudy (abb0t):

That's the same thing, just written differently. But, all you have to do is flip the donimiator and multiply, so instead, you have: \(\sf \large \frac{2}{x^2} \times \color{green}{\frac{x^5}{10}}\)

OpenStudy (anonymous):

@abb0t that's what I did earlier when I tried to solve this, but now would it be cross multiplication like 2*10 and x^2* x^5, or 2*x^5 and x^2*10 ?

OpenStudy (abb0t):

Yes. So do that.

OpenStudy (abb0t):

\(\sf \large \frac{2x^5}{10x^2}\) and reduce.

OpenStudy (anonymous):

thanks ! i just noticed that i made a major mistake at the end when i had solved it cuz instead of going for simplification i divided it which explains why i ended up with a diff answer, but urs is rite. so thank u !i understand it now

OpenStudy (abb0t):

Welcome. Please fan + medal.

OpenStudy (anonymous):

@abb0t could u help me with one more question?